OpenStudy (anonymous):

Please help me with my homework! It is due tomorrow and I don't know how to do this! -.- Solve the system by substitution, make sure to show work. -4.5x-2y=-12.5 3.25x-y=-0.75

jimthompson5910 (jim_thompson5910):

first isolate y in the second equation 3.25x-y=-0.75 -y=-0.75-3.25x -y = -3.25x-0.75 y = 3.25x+0.75

jimthompson5910 (jim_thompson5910):

then plug this into the first equation -4.5x-2y=-12.5 -4.5x-2(3.25x+0.75) = -12.5 now solve for x

OpenStudy (anonymous):

ok, so x=1 now what?

jimthompson5910 (jim_thompson5910):

now plug that into y = 3.25x+0.75 to find y

OpenStudy (anonymous):

y=4

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so x=1 and y=4, is that all that needs to be done?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

thanks :D

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

might have a couple more later